# Distributed High-Dimensional Quantile Regression: Estimation Efficiency and Support Recovery

### v1

@@ 1 Introduction
With the development of modern technology, the proliferation of massive data has garnered significant attention from researchers and practitioners (Li et al., 2020; Gao et al., 2022). For example, financial institutions leverage big data to scrutinize customer preferences and fine-tune their marketing approaches, while manufacturers and transportation departments lean heavily on extensive datasets to streamline supply chain management and enhance delivery route optimization. However, these large-scale datasets are usually distributed arcoss different machines due to storage and privacy concerns, making the direct application of existing statistical methods infeasible. On the other hand, another challenge arises from the high dimensionality of modern data. In the literature, a sparse assumption is often adopted (Zhao & Yu, 2006; Wainwright, 2009; Hastie et al., 2015), and support recovery is an essential problem for high-dimensional analysis. Despite its importance in practice, the support recovery in a distributed system is underexplored in theory, compared to the well-studied statistical estimation of the interested parameters (Zhang et al., 2013; Lee et al., 2017; Jordan et al., 2019).
@@
Since the seminar work of Koenker & Bassett Jr (1978), quantile regression has gained increasing attention across various fields, including economics, biomedicine, and environmental studies. Compared to the least square regression that only estimates the conditional means, quantile regression models the entire conditional quantiles of the response, making it more robust against outliers in the response measurements (Koenker, 2005; Feng et al., 2023). Although quantile regression can handle data heterogeneity better, computational challenges arise when both sample size and dimension are large due to the non-smooth check loss function (He et al., 2023). Consequently, it is natural to consider a distributed estimation procedure to address scalability concerns.
@@
In this paper, we focus on the distributed estimation and support recovery for high-dimensional linear quantile regression. We aim to bridge the theoretical and prcatical gap by addressing several fundamental questions regrading to distributed high-dimensional linear quantile regression. First, what is the statistical limit of estimation in the presence of distributed data? And how does this limit depend on the number of local machines in the distributed system? Second, does distributed high-dimensional linear quantile regression can achieve the same covergence rate of the parameters and support recovery rate as those in a single machine setting? Third, can the proposed method be applied various practical setting, such as homoscedastic and heteroscedastic data structure?
@@
We address the aforementioned theoretical and practical questions by designing a distributed high-dimensional linear quantile regression algorithm via a double smoothing transformation. To our limited knowledge, our work is one of the pioneering works to study the distributed high-dimensional linear quantile regression with the least practical constrain and solid theoretical guarantees involving the estimation efficiency and support recovery. The specific contributions can be concluded as follows:
@@
Methodology novelty. Our paper extends the idea of Chen et al. (2020) resulting in a novel method for Distributed High-dimensional Sparse Quantile Regression (DHSQR) without the stringent assumption that the error term is independent of the covariates. Specifically, we start by transforming the covariate and response, which recast the quantile regression into the least square framework. Next, we introduce an iterative distributed algorithm based on an approximate Newton method that uses a double-smoothing approach applied to the global and local loss functions, respectively. In the distributed system, the local machines only need to broadcast the $(p+1)$-dimensional gradient vectors (instead of the $(p+1)\times(p+1)$ Hessian matrix), where $p$ is the dimension of the covariates. This optimization problem can be efficiently addressed on the central machine due to its simplified least squares formulation.
@@
Theoretical assessments. Theoretically, We not only establish the convergence rate of our DHSQR estimator in $\ell_{2}$-norm (Theorem 3.7 and 3.8), but also characterize the beta-min condition for the exact support recovery (Theorem 3.9 and 3.10) that is novel for distributed high-dimensional sparse estimation (Neykov et al., 2016). After a constant number of iterations, the convergence rate and the beta-min condition of the distributed estimator align with the classical theoretical results derived for a single-machine setup (Wainwright, 2009; Belloni & Chernozhukov, 2011).
@@
Numerical verification. Another contribution of this work is the comprehensive studies on the validity and effectiveness of the proposed algorithm in various synthetic and real-life examples, which further supports the theoretical findings in this paper.
@@ 1.1 Related Work
Distributed methods. Significant efforts have been dedicated to the development of distributed statistical learning methods, broadly categorized into two primary streams. The first class is known as the divide-and-conquer (DC) methods (Zhang et al., 2013; Li et al., 2013; Chen & Xie, 2014; Zhang et al., 2015). These one-shot methods usually compute the relevant estimates based on local samples at the first step and then send these local estimates to a central machine where the final estimate is obtained by simply averaging the local estimates. These methods offer computational efficiency with just one round of communication, but they have the theoretical constraint on the number of local machines to guarantee the global optimal rate (Huang & Huo, 2019; Gao et al., 2022). The second class comprises multi-round distributed methods designed to enhance estimation efficiency and relax restrictions on the number of local machines (Shamir et al., 2014; Lee et al., 2017; Wang et al., 2017). Jordan et al. (2019) and Fan et al. (2021) proposed a communication-efficient surrogate likelihood (CSL) framework that can be applied to low-dimensional estimation, high-dimensional regularized estimation, and Bayesian inference. Notably, the CSL method eliminates the need to transfer local Hessian matrices to the central machine, resulting in significantly reduced communication costs. It’s worth noting that most of the aforementioned methods only focus on homogeneous data, which can be less practical in the context of big data analysis.
@@
Distributed linear quantile regression. In the existing literature, distributed linear quantile regression has been widely investigated using the traditional divide-and-conquer (DC) method (Zhao et al., 2014; Xu et al., 2020; Chen & Zhou, 2020). However, when dealing with high-dimensional settings, where sparsity assumptions are commonly applied, the DC estimator is no longer sparse due to de-biasing and averaging processes, resulting in poor support recovery (Bradic & Kolar, 2017; Chen et al., 2020). In addition, their methods require a limitation on the number of distributed machines to ensure the global convergence rate. To alleviate the restriction on the number of machines, Chen et al. (2020) transform the check loss to the square loss via a kernel smoothing approach and propose a Newton-type distributed estimator. The theoretical guarantees offered insights into estimation errors and support recovery. Nevertheless, their method and theory require the error term to be independent of the covariates which is not very common in practice. Inspired by Jordan et al. (2019) and Fernandes et al. (2021), Tan et al. (2022) studied a distributed convolution-type smoothing quantile regression whose loss function is twice continuously differentiable. We also notice that most of the previous works mainly focus on the convergence rate of their respective estimator, while we further establish the distributed support recovery theory.
@@ 1.2 Notations
[…]
@@ 2 Methodology
In this section, we introduce the proposed distributed estimator. Inspired by the Newton-Raphson iteration, we construct a kernel-based estimator that establishes a connection between quantile regression and ordinary least square regression in a single-machine setting. Based on this estimator, we design a distributed estimation algorithm to minimize a double-smoothing shifted loss function.
@@ 2.1 The Linear Quantile Model
where $Y\in\mathbb{R}$ is a univariate response and ${\boldsymbol{X}}=(x_{0},x_{1},\ldots,x_{p})^{\mathrm{T}}\in\mathbb{R}^{p+1}$ is $p$-dimensional covariate vector with $x_{0}\equiv 1$. Here, $\boldsymbol{\beta}^{*}=\boldsymbol{\beta}^{*}(\tau)=(\beta_{0}^{*}(\tau),\beta_{1}^{*}(\tau),\ldots,\beta_{p}^{*}(\tau))$ is the true coefficient vector that can be obtained by minimizing the following stochastic optimization problem,
@@
where $\rho_{\tau}(u)=u\{\tau-I(u\leq 0)\}$ is the check loss function (Koenker & Bassett Jr, 1978).
@@ 2.2 Newton-type Transformation on Quantile Regression
To solve the stochastic optimization problem (1), we use the Newton-Raphson method. Given an initial estimate $\boldsymbol{\beta}_{0}$, the population form of the Newton-Raphson iteration is
@@
where $\partial\mathcal{Q}(\boldsymbol{\beta})=\boldsymbol{X}\{I(Y-\boldsymbol{X}^{\mathrm{T}}\boldsymbol{\beta}\leq 0)-\tau\}$ is the subgradient of the check loss function with respect to $\boldsymbol{\beta}$, and $\boldsymbol{H}(\boldsymbol{\beta})=\partial\mathbb{E}[\mathcal{Q}(\boldsymbol{\beta})]/\partial\boldsymbol{\beta}=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}}f_{\varepsilon|\boldsymbol{X}}(\boldsymbol{X}^{\mathrm{T}}(\boldsymbol{\beta}-\boldsymbol{\beta}^{*}))]$ denotes the population Hessian matrix of $\mathbb{E}[\mathcal{Q}(\boldsymbol{\beta})]$. Here, $\varepsilon=Y-\boldsymbol{X}^{\mathrm{T}}\boldsymbol{\beta}^{*}$ is the error term and $f_{\varepsilon|\boldsymbol{X}}(x)$ is the conditional density of $\varepsilon$ given $\boldsymbol{X}$.
When the initial estimator $\boldsymbol{\beta}_{0}$ is close to the true parameter $\boldsymbol{\beta}^{*}$, $\boldsymbol{H}(\boldsymbol{\beta}_{0})$ will be close to $\boldsymbol{H}(\boldsymbol{\beta}^{*})=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}}f_{\varepsilon|\boldsymbol{X}}(0)]$. In this paper, we further approximate $\boldsymbol{H}(\boldsymbol{\beta}^{*})$ with $\boldsymbol{D}_{h}(\boldsymbol{\beta}_{0})$ such that
@@
where $e_{0}=Y-{\boldsymbol{X}}^{\mathrm{T}}\boldsymbol{\beta}_{0}$ and $K_{h}(\cdot)=K(\cdot/h)/h$ with $K(\cdot)$ denoting a symmetrix and non-negative kernel function, $h\rightarrow 0$ is the bandwidth. For notation simplicity, we denote $\widetilde{\boldsymbol{X}}_{h}=\sqrt{K_{h}(Y-{\boldsymbol{X}}^{\mathrm{T}}\boldsymbol{\beta})}\boldsymbol{X}$. Replacing $\boldsymbol{H}(\boldsymbol{\beta}_{0})$ with $\boldsymbol{D}_{h}(\boldsymbol{\beta}_{0})$ in (2) leads to the following iteration,
@@
$|\boldsymbol{\beta}_{1}-\boldsymbol{\beta}^{*}|_{2}=\left|\boldsymbol{\beta}_{1}^{*}-\boldsymbol{\beta}^{*}\right|_{2}=O\left(|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}^{*}|_{2}^{2}\right).$
where $\boldsymbol{\beta}_{1}^{*}=\boldsymbol{\beta}_{0}-\boldsymbol{D}_{h}^{-1}(\boldsymbol{\beta}_{0,h})(\boldsymbol{H}(\boldsymbol{\beta}^{*})(\boldsymbol{\beta}_{0}-\boldsymbol{\beta}^{*})+\mathcal{O}(|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}^{*}|_{2}^{2}))$. Consequently, if we have a consistent estimator $\boldsymbol{\beta}_{0}$, we can refine it by (4).
Now we show how to transform the Newton-Raphson iteration into a least square problem. According to (4), we have
@@
If we define a new pseudo response as
@@
then $\boldsymbol{\beta}_{1}=\boldsymbol{D}_{h}^{-1}(\boldsymbol{\beta}_{0})\mathbb{E}(\widetilde{\boldsymbol{X}}_{h}\widetilde{Y}_{h})=\mathop{\rm argmin}_{\boldsymbol{\beta}\in\mathbb{R}^{p+1}}\mathbb{E}(\widetilde{Y}_{h}-\widetilde{\boldsymbol{X}}_{h}^{\mathrm{T}}\boldsymbol{\beta})^{2}$ is the least square regression coefficient of $\widetilde{Y}_{h}$ on $\widetilde{\boldsymbol{X}}_{h}$. To further encourage the sparsity of the coefficient vector, we consider the following $\ell_{1}$-penalized least square problem,
@@
where $\lambda>0$ is the regularization parameter. We can also consider other forms of penalties including the smoothly clipped absolute deviations penalty (SCAD, Fan & Li (2001)) and the minimax concave penalty (MCP, Zhang (2010)). We refer the reader to Hastie et al. (2015) for the comprehensive reviews on the recent developments.
Now we are ready to define the empirical form of $\boldsymbol{\beta}_{1}$ in a single machine. Let $\widehat{\boldsymbol{\beta}}_{0}$ be an initial estimate based on random sample ${\cal Z}^{N}=\{(\boldsymbol{X}_{i},Y_{i})\}_{i=1}^{N}$, then we can transform the origin covariates and responses by
@@
and
$\widetilde{Y}_{i,h}=\widetilde{\boldsymbol{X}}_{i,h}^{\mathrm{T}}\widehat{\boldsymbol{\beta}}_{0}-\frac{1}{\sqrt{K_{h}(e_{0})}}\left(I(\widehat{e}_{0,i}\leq 0)-\tau\right),$ (7)
@@
Given a consistent initial estimator $\widehat{\boldsymbol{\beta}}_{0}$, we introduce a reasonable estimator $\widehat{\boldsymbol{\beta}}_{pool}$ by pooling all the data into a single machine. Compared to the $\ell_{1}$-penalized quantile regression, the least square problem plus a lasso penalty is much more computationally efficient. Moreover, the smoothness and strong convexity of the quadratic loss function facilitate the development of a distributed estimator in the next section.
@@
Different from the work in (Chen et al., 2020), we remove the stringent restriction that the error term $\epsilon$ is independent of the covariate $\boldsymbol{X}$. Therefore we can not simply take the conditional density $f_{\varepsilon|\boldsymbol{X}}(0)$ out of $\boldsymbol{H}(\boldsymbol{\beta}^{*})$ in (3). To consider such a dependence, we further define a pseudo covariate $\widetilde{\boldsymbol{X}}_{h}$ that can be regarded as a density-scaled surrogate of $\boldsymbol{X}$. As indicated in (Chen et al., 2020), the extension to the dependent case seems relatively straightforward in a single-machine setting. It is non-trivial for the distributed implementation both in methodology and theory due to the curse of dimensionality, and we succeed in solving it by leveraging a double-smoothing approach (see the details in next section).

### v3

@@ 1 Introduction
With the development of modern technology, the proliferation of massive data has garnered significant attention from researchers and practitioners (Li et al., 2020; Gao et al., 2022). For example, financial institutions leverage big data to scrutinize customer preferences and fine-tune their marketing approaches, while manufacturers and transportation departments lean heavily on extensive datasets to streamline supply chain management and enhance delivery route optimization. However, these large-scale datasets are usually distributed across different machines due to storage and privacy concerns, making the direct application of existing statistical methods infeasible. On the other hand, another challenge arises from the high dimensionality of modern data. In the literature, a sparse assumption is often adopted (Zhao & Yu, 2006; Wainwright, 2009; Hastie et al., 2015), and support recovery is an essential problem for high-dimensional analysis. Despite its importance in practice, the support recovery in a distributed system is underexplored in theory, compared to the well-studied statistical estimation of the interested parameters (Zhang et al., 2013; Lee et al., 2017; Jordan et al., 2019).
@@
Since the seminar work of Koenker & Bassett Jr (1978), quantile regression has gained increasing attention across various fields, including economics, biomedicine, and environmental studies. Compared to the least squares regression that only estimates the conditional means, quantile regression models the entire conditional quantiles of the response, making it more robust against outliers in the response measurements (Koenker, 2005; Feng et al., 2023). Although quantile regression can better handle data heterogeneity, computational challenges arise when both sample size and dimension are large due to the non-smooth check loss function (He et al., 2023). Consequently, it is natural to consider a distributed estimation procedure to address scalability concerns.
@@
In this paper, we focus on distributed estimation and support recovery for high-dimensional linear quantile regression. We aim to bridge the theoretical and practical gap by addressing several fundamental questions regarding distributed high-dimensional linear quantile regression. First, what is the statistical limit of estimation in the presence of distributed data? And how does this limit depend on the number of local machines in the distributed system? Second, can distributed high-dimensional linear quantile regression achieve the same convergence rate of the parameters and support recovery rate as those in a single machine setting? Third, can the proposed method be applied to various practical settings, such as homoscedastic and heteroscedastic data structures?
@@
We address the aforementioned theoretical and practical questions by designing a distributed high-dimensional linear quantile regression algorithm via a double-smoothing transformation. To our limited knowledge, our work is one of the pioneering works in studying distributed high-dimensional linear quantile regression with the least practical constraint and solid theoretical guarantees involving the estimation efficiency and support recovery. The specific contributions can be concluded as follows.
@@
Methodology novelty. Our paper extends the idea of Chen et al. (2020) resulting in a novel method for Distributed High-dimensional Sparse Quantile Regression (DHSQR) without the stringent assumption that the error term is independent of the covariates. Specifically, we start by transforming the covariate and the response, which recast the quantile regression into the least squares framework. Next, we introduce an iterative distributed algorithm based on an approximate Newton method that uses a double-smoothing approach applied to the global and local loss functions, respectively. In the distributed system, the local machines only need to broadcast the $(p+1)$-dimensional gradient vectors (instead of the $(p+1)\times(p+1)$ Hessian matrix), where $p$ is the dimension of the covariates. This optimization problem can be efficiently addressed on the central machine due to its simplified least squares formulation.
@@
Theoretical assessments. Theoretically, we not only establish the convergence rate of our DHSQR estimator in the $\ell_{2}$-norm (Theorems 3.7 and 3.8), but also characterize the beta-min condition for the exact support recovery (Theorems 3.9 and 3.10) that is novel for distributed high-dimensional sparse estimation (Neykov et al., 2016). After a constant number of iterations, the convergence rate and the beta-min condition of the distributed estimator align with the classical theoretical results derived for a single machine setup (Wainwright, 2009; Belloni & Chernozhukov, 2011).
@@
Numerical verification. Another contribution of this work is the comprehensive studies on the validity and effectiveness of the proposed algorithm in various synthetic and real-life examples, which further support the theoretical findings in this paper.
@@ 1.1 Related Work
Distributed methods. Significant efforts have been dedicated to the development of distributed statistical learning methods, broadly categorized into two main streams. The first class is known as the divide-and-conquer (DC) methods (Zhang et al., 2013; Li et al., 2013; Chen & Xie, 2014; Zhang et al., 2015). These one-shot methods usually compute the relevant estimates based on local samples in the first step and then send these local estimates to a central machine where the final estimate is obtained by simply averaging the local estimates. These methods offer computational efficiency with just one round of communication, but they have the theoretical constraint on the number of local machines to guarantee the global optimal rate (Huang & Huo, 2019; Gao et al., 2022). The second class comprises multi-round distributed methods designed to improve estimation efficiency and relax restrictions on the number of local machines (Shamir et al., 2014; Lee et al., 2017; Wang et al., 2017). Jordan et al. (2019) and Fan et al. (2021) proposed a communication-efficient surrogate likelihood (CSL) framework that can be applied to low-dimensional estimation, high-dimensional regularized estimation, and Bayesian inference. Notably, the CSL method eliminates the need to transfer local Hessian matrices to the central machine, resulting in significantly reduced communication costs. It’s worth noting that most of the aforementioned methods only focus on homogeneous data, which can be less practical in the context of big data analysis.
@@
Distributed linear quantile regression. In the existing literature, distributed linear quantile regression has been widely investigated using the traditional divide-and-conquer method (Zhao et al., 2014; Xu et al., 2020; Chen & Zhou, 2020). However, when dealing with high-dimensional settings, where sparsity assumptions are commonly applied, the DC estimator is no longer sparse due to de-biasing and averaging processes, resulting in poor support recovery (Bradic & Kolar, 2017; Chen et al., 2020). In addition, their methods require a condition on the number of distributed machines to ensure the global convergence rate. To alleviate the restriction on the number of machines, Chen et al. (2020) transform the check loss to the square loss via a kernel smoothing approach and propose a Newton-type distributed estimator. The theoretical results offered insights into estimation errors and support recovery. Nevertheless, their method and theory require the error term to be independent of the covariates, which is not very common and hard to verify in practice. Inspired by the ideas in Jordan et al. (2019) and Fernandes et al. (2021), Tan et al. (2022) studied a distributed convolution-type smoothing quantile regression whose loss function is twice continuously differentiable in both low-dimensional and high-dimensional regime. Note that most of the previous works mainly focus on the convergence rate of their respective estimator, while we further establish the distributed support recovery theory.
@@ 1.2 Notations
For two sequences $\{a_{n}\}$ and $\{b_{n}\}$, we denote $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$, where $C$ is a constant. And $a_{n}\asymp b_{n}$ if and only if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. For a vector $\boldsymbol{u}=(u_{1},\ldots,u_{p})^{\mathrm{T}}$, we define supp$(\boldsymbol{u})=\{j:u_{j}\neq 0\}$. We use $|\cdot|_{q}$ to denote the $\ell_{q}$-norm in $\mathbb{R}_{p}$: $|\boldsymbol{u}|_{q}=(\sum_{i=1}^{p}|u_{i}|^{q})^{1/q}$, for $1\leq q<\infty$ and $|\boldsymbol{u}|_{\infty}=\max_{1\leq i\leq p}|u_{i}|$. For $S\subseteq\{1,\ldots,p\}$ with length $|S|$, let $\boldsymbol{u}_{S}=(u_{i},i\in S)\subseteq\mathbb{R}^{|S|}$. For a matrix $\boldsymbol{A}=(a_{ij})\in\mathbb{R}^{p\times q}$, define $|\boldsymbol{A}|_{\infty}=\max_{1\leq i\leq p,1\leq j\leq q}|a_{ij}|$, $\|\boldsymbol{A}\|_{\infty}=\max_{1\leq i\leq p}\sum_{1\leq j\leq q}|a_{ij}|$, and $\|\boldsymbol{A}\|_{op}=\max_{|\boldsymbol{v}|_{2}=1}|\boldsymbol{A}\boldsymbol{v}|_{2}$. For two subsets $S_{1}\in\{1,\ldots,p\}$ and $S_{2}\in\{1,\ldots,q\}$, we define the submatrix $\boldsymbol{A}_{S_{1}\times S_{2}}=(a_{ij},i\in S_{1},j\in S_{2})$, $\Lambda_{\max}(\boldsymbol{A})$ and $\Lambda_{\min}(\boldsymbol{A})$ to be the largest and smallest eigenvalues of $\boldsymbol{A}$, respectively. In this paper, we use $C$ to denote a universal constant that may vary from line to line.
@@ 2 Methodology
In this section, we introduce the proposed distributed estimator. Inspired by the Newton-Raphson iteration, we construct a kernel-based estimator that establishes a connection between quantile regression and ordinary least squares regression in a single machine setting. Based on this estimator, we design a distributed algorithm to minimize a double-smoothing shifted loss function.
@@ 2.1 The Linear Quantile Model
where $Y\in\mathbb{R}$ is a univariate response and ${\boldsymbol{X}}=(x_{0},x_{1},\ldots,x_{p})^{\mathrm{T}}\in\mathbb{R}^{p+1}$ is $p+1$-dimensional covariate vector with $x_{0}\equiv 1$. Here, $\boldsymbol{\beta}^{*}=\boldsymbol{\beta}^{*}(\tau)=(\beta_{0}^{*}(\tau),\beta_{1}^{*}(\tau),\ldots,\beta_{p}^{*}(\tau))$ is the true coefficient vector that can be obtained by minimizing the following stochastic optimization problem,
@@
where $\rho_{\tau}(u)=u\{\tau-I(u\leq 0)\}$ is the standard check loss function (Koenker & Bassett Jr, 1978) with $I(\cdot)$ is the indicator function.
@@ 2.2 Newton-type Transformation on Quantile Regression
To solve the stochastic optimization problem in (1), we use the Newton-Raphson method. Given an initial estimator $\boldsymbol{\beta}_{0}$, the population form of the Newton-Raphson iteration is
@@
where $\partial\mathcal{Q}(\boldsymbol{\beta})=\boldsymbol{X}\{I(Y-\boldsymbol{X}^{\mathrm{T}}\boldsymbol{\beta}\leq 0)-\tau\}$ is the subgradient of the check loss function with respect to $\boldsymbol{\beta}$, and $\boldsymbol{H}(\boldsymbol{\beta})=\partial\mathbb{E}[\partial\mathcal{Q}(\boldsymbol{\beta})]/\partial\boldsymbol{\beta}=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}}f_{\varepsilon|\boldsymbol{X}}(\boldsymbol{X}^{\mathrm{T}}(\boldsymbol{\beta}-\boldsymbol{\beta}^{*}))]$ denotes the population Hessian matrix of $\mathbb{E}[\mathcal{Q}(\boldsymbol{\beta})]$. Here, we denote the error term as $\varepsilon=Y-\boldsymbol{X}^{\mathrm{T}}\boldsymbol{\beta}^{*}$, and $f_{\varepsilon|\boldsymbol{X}}(\cdot)$ is the conditional density of $\varepsilon$ given $\boldsymbol{X}$.
When the initial estimator $\boldsymbol{\beta}_{0}$ is close to the true parameter $\boldsymbol{\beta}^{*}$, $\boldsymbol{H}(\boldsymbol{\beta}_{0})$ will be close to $\boldsymbol{H}(\boldsymbol{\beta}^{*})=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{\mathrm{T}}f_{\varepsilon|\boldsymbol{X}}(0)]$. Motivated by this, we further approximate $\boldsymbol{H}(\boldsymbol{\beta}^{*})$ with $\boldsymbol{D}_{h}(\boldsymbol{\beta}_{0})$ such that
@@
where $e_{0}=Y-{\boldsymbol{X}}^{\mathrm{T}}\boldsymbol{\beta}_{0}$, and $K_{h}(\cdot)=K(\cdot/h)/h$ with $K(\cdot)$ denoting a symmetrix and non-negative kernel function, $h\rightarrow 0$ is the bandwidth. For simplicity of notation, we denote a pseudo covariate as $\widetilde{\boldsymbol{X}}_{h}=\sqrt{K_{h}(e_{0})}\boldsymbol{X}$. Hence, we can rewrite $\boldsymbol{D}_{h}(\boldsymbol{\beta}_{0})=\mathbb{E}(\widetilde{\boldsymbol{X}}_{h}\widetilde{\boldsymbol{X}}_{h}^{\mathrm{T}})$, which is the covariance matrix of $\widetilde{\boldsymbol{X}}_{h}$. Replacing $\boldsymbol{H}(\boldsymbol{\beta}_{0})$ with $\boldsymbol{D}_{h}(\boldsymbol{\beta}_{0})$ in (2) leads to the following iteration,
@@
$|\boldsymbol{\beta}_{1}-\boldsymbol{\beta}^{*}|_{2}$ $=\left|\boldsymbol{\beta}_{0}-\boldsymbol{D}_{h}^{-1}(\boldsymbol{\beta}_{0,h})(\boldsymbol{H}(\boldsymbol{\beta}^{*})(\boldsymbol{\beta}_{0}-\boldsymbol{\beta}^{*})\right.$
$\left.+\mathcal{O}(|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}^{*}|_{2}^{2}))-\boldsymbol{\beta}^{*}\right|_{2}=\mathcal{O}\left(|\boldsymbol{\beta}_{0}-\boldsymbol{\beta}^{*}|_{2}^{2}\right).$
Consequently, if we have a consistent estimator $\boldsymbol{\beta}_{0}$, we can refine it by (4).
Now we show how to transform the Newton-Raphson iteration into a least squares problem. According to (4), we have
@@
If we further define a new pseudo response as
@@
then $\boldsymbol{\beta}_{1}=\boldsymbol{D}_{h}^{-1}(\boldsymbol{\beta}_{0})\mathbb{E}(\widetilde{\boldsymbol{X}}_{h}\widetilde{Y}_{h})=\mathop{\rm argmin}_{\boldsymbol{\beta}\in\mathbb{R}^{p+1}}\mathbb{E}(\widetilde{Y}_{h}-\widetilde{\boldsymbol{X}}_{h}^{\mathrm{T}}\boldsymbol{\beta})^{2}$ is the least squares regression coefficient of $\widetilde{Y}_{h}$ on $\widetilde{\boldsymbol{X}}_{h}$. To further encourage the sparsity of the coefficient vector, we consider the following $\ell_{1}$-penalized least squares problem,
@@
where $\lambda>0$ is the regularization parameter. We can also consider other forms of penalties including the smoothly clipped absolute deviations penalty (SCAD, Fan & Li (2001)) and the minimax concave penalty (MCP, Zhang (2010)). We refer the reader to Hastie et al. (2015) for comprehensive reviews on the recent developments.
Now we are ready to define the empirical form of $\boldsymbol{\beta}_{1}$ in a single machine. Let $\widehat{\boldsymbol{\beta}}_{0}$ be an initial estimate based on random samples ${\cal Z}^{N}=\{(\boldsymbol{X}_{i},Y_{i})\}_{i=1}^{N}$, then we can transform the origin covariates and responses by
@@
$\widetilde{Y}_{i,h}=\widetilde{\boldsymbol{X}}_{i,h}^{\mathrm{T}}\widehat{\boldsymbol{\beta}}_{0}-\frac{1}{\sqrt{K_{h}(\widehat{e}_{0,i})}}\left(I(\widehat{e}_{0,i}\leq 0)-\tau\right),$ (7)
@@
Given a consistent initial estimator $\widehat{\boldsymbol{\beta}}_{0}$, we introduce a reasonable estimator $\widehat{\boldsymbol{\beta}}_{pool}$ by pooling all data into a single machine. Compared to the standard $\ell_{1}$-penalized quantile regression, the least squares problem plus a lasso penalty is much more computationally efficient. Moreover, the smoothness and strong convexity of the quadratic loss function facilitate the development of a distributed estimator in the next section.
@@
Different from the work in Chen et al. (2020), we remove the stringent restriction that the error term $\epsilon$ should be independent of the covariate $\boldsymbol{X}$. Therefore, we cannot simply take the conditional density $f_{\varepsilon|\boldsymbol{X}}(0)$ from $\boldsymbol{H}(\boldsymbol{\beta}^{*})$ in (3). To consider such a dependence, we further define a pseudo covariate $\widetilde{\boldsymbol{X}}_{h}$ that can be regarded as a density-scaled surrogate of $\boldsymbol{X}$. As indicated in Chen et al. (2020), the extension to the dependent case seems relatively straightforward in a single machine setting. However, it is nontrivial for distributed implementation both in methodology and theory due to the curse of dimensionality, and we succeed in solving it by leveraging a double-smoothing approach (see details in the next section).